# Minimizing Human Intervention in Online Classification

### v1

@@ Abstract
We introduce and study an online problem arising in question answering systems. In this problem, an agent must sequentially classify user-submitted queries represented by $d$-dimensional embeddings drawn i.i.d. from an unknown distribution. The agent may consult a costly human expert for the correct label, or guess on her own without receiving feedback. The goal is to minimize regret against an oracle with free expert access. When the time horizon $T$ is at least exponential in the embedding dimension $d$, one can learn the geometry of the class regions: in this regime, we propose the Conservative Hull-based Classifier (CHC), which maintains convex hulls of expert-labeled queries and calls the expert as soon as a query lands outside all known hulls. CHC attains $\mathcal{O}(\log^{d}T)$ regret in $T$ and is minimax optimal for $d=1$. Otherwise, the geometry cannot be reliably learned without additional distributional assumptions. We show that when the queries are drawn from a subgaussian mixture, for $T\leq e^{d}$, a Center-based Classifier (CC) achieves regret proportional to $N\log{N}$ where $N$ is the number of labels. To bridge these regimes, we introduce the Generalized Hull-based Classifier (GHC), a practical extension of CHC that allows for more aggressive guessing via a tunable threshold parameter. Our approach is validated with experiments, notably on real-world question-answering datasets using embeddings derived from state-of-the-art large language models.
1 INTRODUCTION
@@
Question answering and customer support systems often start with no prior knowledge and must learn to respond to user queries over time. This work introduces an online classification problem addressing this challenge. Initially, human experts are needed to answer incoming queries, and each answered query is stored to aid future responses. However, expert interventions are costly, and the system should minimize unnecessary queries to the expert. This raises a fundamental question: How can we design a system that progressively builds competence in answering queries while minimizing reliance on expert interventions? We address this question in a setting where each query is represented by an embedding, typically produced by a language model. We formalize the problem as an online decision-making task, that we call online classification with expert guidance. At each round $t=1,2,\dots$, an agent observes a query embedding $q_{t}\in\mathbb{R}^{d}$ drawn i.i.d. from an unknown distribution. There are $N$ latent answer labels, each corresponding to some unknown region in the embedding space. The agent must assign a label to the query. She may either predict a label immediately or ask an expert to obtain the true label (incurring a cost). Crucially, if the agent guesses on its own, no label is revealed. The goal is to devise algorithms that minimize the cumulative cost after $T$ queries, where the costs come from wrongly guessing the label or calling the expert.
Besides the introduction of the online classification with expert guidance problem, our contributions are as follows:
@@
a. Algorithms We introduce two main algorithms: the Conservative Hull-based Classifier (CHC) and its generalization, GHC. Both exploit the geometry of class regions in the embedding space to guide decisions. CHC is conservative and never makes incorrect predictions; GHC introduces a tunable threshold that allows for more frequent guessing. This threshold controls the trade-off between accuracy and expert use, and is particularly beneficial in high-dimensional spaces such as those defined by LLM embeddings.
@@
b. Theoretical Guarantees We provide the first regret bounds for the online classification with expert guidance problem. Using tools from convex geometry, we show that CHC achieves a regret scaling as $\mathcal{O}(\log^{d}{T})$ after $T$ queries, where $d$ is the embedding dimension. These bounds are distribution-free and robust. Moreover, we show that in dimension $d=1$, CHC is minimax optimal. To complement our distribution-free bounds, we additionally show that for well-separated subgaussian mixtures with equal weights, a Center-based Classifier (CC) attains regret proportional to $N\log{N}$ when $T\leq e^{d}.$
@@
c. Empirical Validation We evaluate CHC, CC and GHC on both synthetic data and real-world question-answering tasks, using embeddings from various Large Language Models (LLMs). We demonstrate that leveraging higher-dimensional embeddings significantly reduces the overall regret. Additionally, we show that GHC outperforms CC and a sequential $k$-means baseline in all settings.
@@ 2 RELATED WORK
In this section, we review the literature on online classification and clustering. Additional related work on embedding-based retrieval and convex geometry is presented in Appendix D.
Online classification has a rich history in theoretical computer science and statistical machine learning (Littlestone,, 1987; Ben-David et al.,, 2009; Rakhlin et al.,, 2015). A recent line of work studies streaming classification, where a learner must predict binary labels (Ben-Eliezer and Yogev,, 2020; Alon et al.,, 2021; Montasser et al.,, 2025). These works differ from ours in several key respects: (i) their setting is adversarial; (ii) when the learner predicts incorrectly, the correct label is revealed; and (iii) their models do not involve the type of decision-making trade-offs present in ours.
In online classification with limited label information, Cesa-bianchi et al., (2004) studied binary online classification where the learner actively decides whether to ask for the correct label. Cesa-Bianchi et al., (2005) similarly considered a feedback model where information is only gained by an explicit action. Their model is adversarial, and less information is available to the learner than in our model: only the labels are observed. They obtain sublinear regret, but compared to a much easier baseline than ours, namely the best algorithm that guesses the same label at each round, whereas we compare to an oracle that always guesses the correct label. Bechavod et al., (2019) further studied one-sided feedback, revealing labels only upon positive predictions in an i.i.d. stream. In this work, we specifically exploit the geometric structure of the class regions to decide when to call the expert.
Finally, Kane et al., (2017) consider a pool-based active learning model, where the entire dataset is available from the start. Their goal is exact recovery of all labels for the entire pool, and they allow comparison queries ("which point is closer to the boundary between classes?"). In comparison, our model operates in an i.i.d. online streaming setting, and our notion of regret aims at balancing the costs of wrong guesses and expert calls.
@@
Clustering refers to the task of grouping data points based on similarity, and is hence related to classification. Clustering is widely used to summarize and organize large datasets (Pandove et al.,, 2018), typically under the assumption that all data is available upfront in a batch setting. Clustering in high dimensions is notoriously difficult, both statistically and computationally (Steinbach et al.,, 2004). To overcome this, most batch clustering theoretical works assume strong generative or separability conditions. A rich line of work aims at characterizing the optimal clustering error rate and devising algorithms reaching these limits. For isotropic Gaussian Mixture Models (GMMs) (Pearson,, 1894; Renshaw et al.,, 1987), it is shown in Lu and Zhou, (2016) that the average number of misclustered points of Lloyd’s algorithm (Lloyd,, 1982) achieves the minimax error rate. More recent analyses extend to slightly more general settings, for example anisotropic GMMs (Chen and Zhang,, 2024) and mixture with sub-exponential tails (Dreveton et al.,, 2024), for which optimal rates are obtained with a variant of Lloyd’s algorithm. All such results rely on strong distributional assumptions. In contrast, our framework is distribution-free, and we notably derive regret bounds that hold even in the uniform setting, where no natural clusters exist. In fact, the class regions in our setting need only be defined by convex polytopes of the embedding space, a structure we use to specify the expert’s labeling policy. Furthermore, our framework is online and incorporates a decision trade-off between prediction and expert consultation not found in these works.
@@
Many modern applications—such as news feeds, social media, and interactive search—generate data as continuous streams. These scenarios have motivated the development of online clustering methods, which must operate in real time as data arrives. Choromanska and Monteleoni, (2012) propose an online clustering algorithm that maintains an ensemble of batch $k$-means algorithms as “experts” and re-weights them by approximating their current clustering cost. Cohen-Addad et al., (2021) consider online $k$-means clustering: at each step the algorithm maintains $k$ centers and incurs loss equal to the squared distance from the new point to the nearest center, aiming to minimize regret against the best fixed clustering in hindsight. These methods focus on minimizing a distance-based cost rather than the classification/regret perspective we consider. Moreover, they do not capture the fundamental trade-off in our model between querying an expert and making an unsupported prediction.
@@ 3.1 Online Classification Problem
a. Query A user generates a query characterized by its embedding or representation $q_{t}\in{\cal E}\subset\mathbb{R}^{d}$ in an i.i.d. manner according to an unknown distribution $\mu$. The agent must address the query. We assume that there are $N$ possible labels $i\in[N]:=\{1,\ldots,N\}$. Throughout the paper, the embedding space is either the hypercube ${\cal E}={\cal I}^{d}:=[0,1]^{d}$ or the unit sphere ${\cal E}=\mathcal{S}^{d-1}:=\{x\in\mathbb{R}^{d}:\|x\|_{2}=1\}$ (this is usually the case when the embedding comes from an LLM).
@@ 3.2 Voronoi Regret
A natural example of partition is constructed as follows. Assume that each label $i$ also has a representation $s_{i}\in{\cal E}$ that is referred to as its seed. $s_{i}$ may be interpreted as the typical query corresponding to the correct label $i$ (ii is the correct label to queries whose embeddings would be close to $s_{i}$). $\mathcal{E}$ can then be partitioned as the Voronoi tessellation generated by the seeds $s_{1},\ldots,s_{N}$. This means that for any $t$, $i_{t}\in\arg\min_{i\in[N]}\|q_{t}-s_{i}\|_{2}$ and for any $i$, $\mathcal{C}_{i}=\{q\in\mathcal{E}:\|q-s_{i}\|_{2}\leq\|q-s_{j}\|_{2},\;\forall j\in[N]\setminus\{i\}\}$. This example of partition is justified as follows: When the query distribution $\mu$ is such that $\mathbb{P}(i_{t}=i)=1/N$ and has a conditional density $p(q_{t}|i_{t}=i)=f(\|s_{i}-q_{t}\|_{2})$ where f is decreasing, the expert’s label corresponds to the Maximum Likelihood Estimator (MLE) for the correct label, assuming the $s_{i}$ are known. Furthermore, when $\mathcal{E}=\mathcal{S}^{d-1}$ is the unit sphere, the expert assigns to a query $q$ the index of the seed maximizing the cosine similarity $q^{\top}s_{i}$, as $q,s_{i}$ have unit norm (cosine similarity is a standard measure of semantic similarity in embedding-based Natural Language Processing (NLP) models (Reimers and Gurevych,, 2019)). In this case, the boundaries of the Voronoi cells are arcs of great circles, and the resulting cells form spherical convex polytopes, as illustrated in Fig. 1.
Since the Oracle algorithm knows the correct labels, it earns a reward $\beta$ per round. We hence define the regret of an algorithm $\pi$ up to round $T$ as: $R_{\pi}(T)=\beta T-\sum_{t=1}^{T}\mathbb{E}[r_{\pi}(t)]$. Our objective is to devise an algorithm that minimizes this regret.
@@ 3.3 Notation and Assumptions
When ${\cal E}=\mathcal{S}^{d-1}$, the assumption that each cell lies within an open half-sphere is instrumental for applying the gnomonic projection (see Section 3.1 in Besau and Schuster, (2016)) in our regret analysis. Let $V$ denote the natural volume measure on $\mathcal{E}$, namely the Lebesgue measure $\lambda$ on $\mathbb{R}^{d}$ if $\mathcal{E}=\mathcal{I}^{d}$, or the spherical Lebesgue measure $\omega$ on $\mathcal{S}^{d-1}$ if $\mathcal{E}=\mathcal{S}^{d-1}$.
@@ 4.1 Conservative Hull-Based Classifier (CHC)
Importantly, to implement CHC, we do not need to compute convex hulls (this would be computationally prohibitive—algorithms such as QuickHull have a worst-case complexity that is exponential in $d/2$, see Theorem 3.2 in Barber et al., (1996)). Indeed, we only need to be able to check if a query $q\in\mathcal{E}$ belongs to the convex hull of some finite set of queries ${\cal Q}_{i}=\{p_{1},\dots,p_{n}\}$ . If $\mathcal{E}=\mathcal{I}^{d}$, this amounts to checking if there exists $\alpha\geq 0\in\mathbb{R}^{n}$ such that $\sum_{i=1}^{n}\alpha_{i}=1$ and $q=\sum_{i=1}^{n}\alpha_{i}p_{i}$. If $\mathcal{E}=\mathcal{S}^{d-1}$, the same condition must be checked but without requiring $\sum_{i=1}^{n}\alpha_{i}=1$. Checking these conditions can be formulated as linear programs with $n$ variables and $n+d+1$ or $n+d$ constraints respectively. These programs can be solved in a time polynomial in the dimension and the number of queries by interior-point methods, see e.g. Khachiyan, (1979); Karmarkar, (1984).
@@
To state the regret guarantees for CHC, we need to define the number $F(P)$ of flags of a polytope $P$. A flag of $P$ is a sequence $(F_{j})_{j=0}^{d-1}$ of faces [^1] Faces of a polytope are specific planar surfaces on its boundary, e.g., for a cube, the 0-dim faces are the vertices, the $1$-dim faces are the $12$ edges, and the $2$-dim faces are the $6$ squares forming the boundary. of $P$ such that $\dim(F_{j})=j$ and $F_{0}\subset\dots\subset F_{d-1}$. For example, for $d=2$ and $P=[0,1]^{2}$, a flag is a sequence formed by a vertex of the square and an edge incident to that vertex, and $F(P)=4\times 2=8$. Refer to Section 2.1 in Besau et al., (2018) for details. The following regret guarantees for CHC hold for any convex polytope partition $\{\mathcal{C}_{i}\}_{i\in[N]}$ of the embedding space by the expert.
@@
To upper bound $\mathbb{E}[\mu(\mathcal{C}_{i}\setminus\hat{\mathcal{C}}_{i,t})]$, we build upon a result from Bárány and Buchta, (1993), which characterizes the expected volume of the convex hull of $n$ points sampled uniformly at random from a convex polytope in $\mathbb{R}^{d}$. We extend this result in two directions: (i) to distributions with bounded density via a rejection sampling argument; (ii) to the spherical setting where points lie on $\mathcal{S}^{d-1}$, by mapping spherical polytopes to Euclidean polytopes in $\mathbb{R}^{d-1}$ using the gnomonic projection. The dependence on $C/c$ (and $CK/c$ in the spherical case) arises from the rejection sampling step and is likely an artifact of the proof technique—we conjecture that this dependence could be removed with a tighter analysis.
@@
Voronoi Regret. We finally derive explicit regret upper bounds when the cells form a Voronoi tessellation, as described in Section 3.2. The following corollary is obtained from Theorem 4.1 by controlling the total number of flags, $\sum_{i=1}^{N}F(\mathcal{C}_{i})$. The proof, given in Appendix B.1.2, leverages an upper bound on the number of faces of a convex polytope from McMullen, (1970), along with the observation that the number of facets of each cell $\mathcal{C}_{i}$ is at most N1+2dN-1+2d when $\mathcal{E}={\cal I}^{d}$ (the hypercube), and at most N1N-1 when $\mathcal{E}={\cal S}^{d-1}$ (the unit sphere).
@@
It may seem counterintuitive that the constant preceding the leading term (e.g., $\log^{d}T$ when ${\cal E}={\cal I}^{d}$) decreases with increasing $d$. This behavior suggests that the error term conceals a more intricate dependency on the dimension. Precisely characterizing this dependency is difficult and touches upon unresolved questions in convex geometry (see Appendix E.3 for further discussion). Nonetheless, our upper bounds on the regret of CHC are most informative in regimes where $T$ exceeds a threshold that grows exponentially with $d$, or equivalently, when the dimension is not excessively large. This is consistent with results from Chakraborti et al., (2021), which show that the expected volume of a convex hull formed by random points remains negligible until the number of samples is exponential in $d$. In our setting, this implies that a significant reduction in expert queries only begins once this sample threshold is crossed.

### v2

@@ Abstract
Training or fine-tuning large language model (LLM)–based systems often requires costly human feedback, yet there is limited understanding of how to minimize such intervention while maintaining strong error guarantees. We study this problem for LLM-based classification systems in an active learning framework: an agent sequentially labels $d$-dimensional query embeddings drawn i.i.d. from an unknown distribution by either calling a costly expert or guessing with no feedback, with the goal of minimizing regret relative to an oracle with free expert access. When the horizon $T$ is at least exponential in the embedding dimension $d$, the geometry of the class regions can be learned. In this regime, we propose the Conservative Hull-based Classifier (CHC), which maintains convex hulls of expert-labeled queries and calls the expert when a query lands outside all known hulls. CHC attains $\mathcal{O}(\log^{d}T)$ regret in $T$ and is minimax optimal for $d=1$. Otherwise, the geometry cannot be reliably learned in general. We show that for queries drawn from a subgaussian mixture and $T\leq e^{d}$, a Center-based Classifier (CC) achieves regret proportional to $N\log{N}$ where $N$ is the number of labels. To bridge these regimes, we introduce the Generalized Hull-based Classifier (GHC), a practical extension of CHC that enables more aggressive guessing via a tunable parameter. Our approach is validated on real-world question-answering datasets using state-of-the-art text embedding models.
@@
Question answering and customer support systems often start with no prior knowledge and must learn to respond to user queries over time. Initially, human experts are needed to answer incoming queries, and each answered query is stored to aid future responses. However, expert interventions are costly, and the system should minimize unnecessary queries to the expert. This raises a fundamental question: How can we design a system that progressively builds competence in answering queries while minimizing reliance on expert interventions? We address this question in a setting where each query is represented by an embedding, typically produced by a language model. We formalize this problem as an active learning task. At each round $t=1,2,\dots$, an agent observes a query embedding $q_{t}\in\mathbb{R}^{d}$ drawn i.i.d. from an unknown distribution. There are $N$ latent answer labels, each corresponding to some unknown region in the embedding space. The agent must assign a label to the query: she may either predict a label immediately or ask an expert to obtain the true label (incurring a cost). Crucially, if the agent guesses on its own, no label is revealed. The goal is to devise algorithms that minimize the cumulative cost after $T$ queries, where the cost comes from wrongly guessing the label or calling the expert. Our contributions are as follows:
@@
a. Algorithms We introduce three algorithms: most importantly, the Conservative Hull-based Classifier (CHC) and its generalization, GHC. Both exploit the geometry of class regions in the embedding space to guide decisions. CHC is conservative and never makes incorrect predictions; GHC introduces a tunable threshold that allows for more frequent guessing. This threshold controls the trade-off between accuracy and expert use, and is particularly beneficial in high-dimensional spaces such as those defined by LLM embeddings. We additionally present a simpler Center-Based Classifier (CC) that relies on estimates of the class regions centers to make guesses.
@@
b. Theoretical Guarantees We provide bounds on the regret of CHC and CC. Using tools from convex geometry, we show that CHC achieves a regret scaling as $\mathcal{O}(\log^{d}{T})$ after $T$ queries, where $d$ is the embedding dimension. These bounds are distribution-free and robust. Moreover, we show that in dimension $d=1$, CHC is minimax optimal. To complement our distribution-free bounds, we additionally show that for well-separated subgaussian mixtures with equal weights, CCattains regret proportional to $N\log{N}$ when $T\leq e^{d}.$
@@
c. Empirical Validation We evaluate CHC, CC and GHC on both synthetic data and real-world question-answering tasks, using embeddings from various text embedding models. We demonstrate that leveraging higher-dimensional embeddings significantly reduces the overall regret. Additionally, we show that GHC outperforms CC and baselines from the literature in all settings.
@@ 1 RELATED WORK
There is a growing literature on learning problems where an agent must decide when to defer to an expert or abstain from providing an answer. Our work is related to two frameworks where this objective is central: learning to defer and active learning. We further review adjacent literature on clustering. Additional related work on online classification, embedding-based retrieval and convex geometry is presented in Appendix D.
Learning to Defer studies how to combine automated predictions with human oversight (Madras et al., 2018; Mozannar and Sontag, 2020). It is framed as an offline problem where the learner has access to a training dataset of inputs, ground-truth labels, and expert predictions. The primary objective is to minimize a test-time loss that accounts for both prediction error and the cost of human intervention. Our online setting differs fundamentally: rather than leveraging a pre-existing dataset, the agent faces a cold-start problem and must decide at each round whether to call the expert, building competence over time. Consequently, we evaluate performance via cumulative regret rather than test-time risk.
Active learning. Online learning has a rich history in theoretical computer science and statistical machine learning (Littlestone, 1987; Ben-David et al., 2009; Rakhlin et al., 2015). Of particular relevance to our work is the subfield of active learning, where the learner decides when to query for a label to minimize annotation costs (Dasgupta, 2011; Hanneke, 2014). Much of the literature addresses the pool-based setting, where the entire dataset is available from the start (Lewis and Gale, 1994; Houlsby et al., 2011; Kane et al., 2017; Gentile et al., 2022), and the goal is to minimize the number of queries required to learn a classifier with low error rate, ignoring prediction errors during training. In the stream-based setting, data points are observed sequentially, but the objective typically remains minimizing the error rate at test-time instead of the cumulative number of prediction errors (Freund et al., 1997; Sabato and Hess, 2018). A foundational approach to this problem is the CAL algorithm (Cohn et al., 1994), which queries for the label as soon as there is some disagreement among classifiers consistent with previously queried labels, i.e., it queries only when the label cannot be inferred. Although known to be optimal in terms of the number of label requests among zero-error classifiers in the realizable setting (Hanneke, 2014), it is generally computationally inefficient. Our CHC algorithm (Section 3.1) can be understood as a computationally efficient approximation of CAL, when class membership is dictated by a convex partition of the input space. Our setting is most similar to online selective sampling (Hanneke and Yang, 2021), where the object of study is the same as ours, i.e., the trade-off between the number of prediction errors and the number of label queries. Several theoretical works focus on linear binary classification with inputs either adversarially selected with a hard margin constraint (Cesa-bianchi et al., 2004; Lu et al., 2016), or stochastic under the Tsybakov noise condition (Wang and Singh, 2016), whereas we study the multiclass setting and derive bounds over a wide class of distributions by exploiting the geometric structure of the class regions.
Online Selective Classification. Finally, our setting is closely related to selective classification (also known as classification with a reject option), where the learner may opt to not make a prediction. Typically, data is available in batches (Chow, 1970; Bartlett and Wegkamp, 2008; El-Yaniv and Wiener, 2010), and the goal is to balance the prediction error probability assuming a prediction is made with the probability of making a prediction, also known as the coverage. This setting was shown to bear similarity with stream-based active learning in El-Yaniv and Wiener (2012). Recent work has explored online variants where labels are revealed only when the learner predicts (Cortes et al., 2018; Bechavod et al., 2019) or abstains (Gangrade et al., 2021). The latter work is the most relevant to ours. They derive bounds on the cumulative number of incorrect guesses and label queries, but assuming a finite set of candidate classifiers, whereas ours is infinite (all classifiers partitioning the embedding space as convex regions).
@@
Clustering refers to the task of grouping data points based on similarity, and is hence related to our problem. It is widely used to summarize and organize large datasets (Pandove et al., 2018), typically under the assumption that all data is available upfront in a batch setting. Clustering in high dimensions is notoriously difficult, both statistically and computationally (Steinbach et al., 2004). To overcome this, most batch clustering theoretical works assume strong generative or separability conditions. A rich line of work aims at characterizing the optimal clustering error rate and devising algorithms reaching these limits. For isotropic Gaussian Mixture Models (GMMs) (Pearson, 1894; Renshaw et al., 1987), it is shown in Lu and Zhou (2016) that the average number of misclustered points of Lloyd’s algorithm (Lloyd, 1982) achieves the minimax error rate. More recent analyses extend to slightly more general settings, for example anisotropic GMMs (Chen and Zhang, 2024) and mixture with sub-exponential tails (Dreveton et al., 2024), for which optimal rates are obtained with a variant of Lloyd’s algorithm. All such results rely on strong distributional assumptions. In contrast, our framework is distribution-free, and we notably derive regret bounds that hold even in the uniform setting, where no natural clusters exist. In fact, the class regions in our setting need only be defined by convex polytopes of the embedding space, a structure we use to specify the expert’s labeling policy. Furthermore, our framework is online and incorporates a decision trade-off between prediction and expert consultation not found in these works.
@@
Many modern applications—such as news feeds, social media, and interactive search—generate data as continuous streams. These scenarios have motivated the development of online clustering methods, which must operate in real time as data arrives. Choromanska and Monteleoni (2012) propose an online clustering algorithm that maintains an ensemble of batch $k$-means algorithms as “experts” and re-weights them by approximating their current clustering cost. Cohen-Addad et al. (2021) consider online $k$-means clustering: at each step the algorithm maintains $k$ centers and incurs loss equal to the squared distance from the new point to the nearest center, aiming to minimize regret against the best fixed clustering in hindsight. These methods focus on minimizing a distance-based cost rather than the classification/regret perspective we consider. Moreover, they do not capture the fundamental trade-off in our model between querying an expert and making an unsupported prediction.
@@ 2.1 Online Classification Problem
a. Query A user generates a query characterized by its embedding or representation $q_{t}\in{\cal E}\subset\mathbb{R}^{d}$ in an i.i.d. manner according to an unknown distribution $\mu$. The agent must address the query. We assume that there are $N$ possible labels $i\in[N]:=\{1,\ldots,N\}$. Throughout the paper, the embedding space is either the hypercube ${\cal E}={\cal I}^{d}:=[0,1]^{d}$ or the unit sphere ${\cal E}=\mathcal{S}^{d-1}:=\{x\in\mathbb{R}^{d}:\|x\|_{2}=1\}$ (this is usually the case with text embedding models).
@@ 2.2 Voronoi Regret
A natural example of partition is constructed as follows. Assume that each label $i$ also has a representation $s_{i}\in{\cal E}$ that is referred to as its seed. $s_{i}$ may be interpreted as the typical query corresponding to the correct label $i$ (ii is the correct label to queries whose embeddings would be close to $s_{i}$). $\mathcal{E}$ can then be partitioned as the Voronoi tessellation generated by the seeds $s_{1},\ldots,s_{N}$. This means that for any $t$, $i_{t}\in\arg\min_{i\in[N]}\|q_{t}-s_{i}\|_{2}$ and for any $i$, $\mathcal{C}_{i}=\{q\in\mathcal{E}:\|q-s_{i}\|_{2}\leq\|q-s_{j}\|_{2},\;\forall j\in[N]\setminus\{i\}\}.$ This example of partition is justified as follows: When the query distribution $\mu$ is such that $\mathbb{P}(i_{t}=i)=1/N$ and has a conditional density $p(q_{t}|i_{t}=i)=f(\|s_{i}-q_{t}\|_{2})$ where f is decreasing, the expert’s label corresponds to the Maximum Likelihood Estimator (MLE) for the correct label, assuming the $s_{i}$ are known. Furthermore, when $\mathcal{E}=\mathcal{S}^{d-1}$ is the unit sphere, the expert assigns to a query $q$ the index of the seed maximizing the cosine similarity $q^{\top}s_{i}$, as $q,s_{i}$ have unit norm (cosine similarity is a standard measure of semantic similarity in embedding-based Natural Language Processing (NLP) models (Reimers and Gurevych, 2019)). In this case, the boundaries of the Voronoi cells are arcs of great circles, and the resulting cells form spherical convex polytopes, as illustrated in Fig. 1.
Since the Oracle algorithm knows the correct labels, it earns a reward $\beta$ per round. We hence define the regret of an algorithm $\pi$ up to round $T$ as
$R_{\pi}(T)=\beta T-\sum_{t=1}^{T}\mathbb{E}[r_{\pi}(t)].$
Our objective is to devise an algorithm that minimizes this regret.
@@ 2.3 Notation and Assumptions
When ${\cal E}=\mathcal{S}^{d-1}$, the assumption that each cell lies within an open half-sphere is instrumental for applying the gnomonic projection (see Section 3.1 in Besau and Schuster (2016)) in our regret analysis. Let $V$ denote the natural volume measure on $\mathcal{E}$, namely the Lebesgue measure $\lambda$ on $\mathbb{R}^{d}$ if $\mathcal{E}=\mathcal{I}^{d}$, or the spherical Lebesgue measure $\omega$ on $\mathcal{S}^{d-1}$ if $\mathcal{E}=\mathcal{S}^{d-1}$.
@@ 3.1 Conservative Hull-Based Classifier (CHC)
Importantly, to implement CHC, we do not need to compute convex hulls (this would be computationally prohibitive—algorithms such as QuickHull have a worst-case complexity that is exponential in $d/2$, see Theorem 3.2 in Barber et al. (1996)). Indeed, we only need to be able to check if a query $q\in\mathcal{E}$ belongs to the convex hull of some finite set of queries ${\cal Q}_{i}=\{p_{1},\dots,p_{n}\}$ . If $\mathcal{E}=\mathcal{I}^{d}$, this amounts to checking if there exists $\alpha\geq 0\in\mathbb{R}^{n}$ such that $\sum_{i=1}^{n}\alpha_{i}=1$ and $q=\sum_{i=1}^{n}\alpha_{i}p_{i}$. If $\mathcal{E}=\mathcal{S}^{d-1}$, the same condition must be checked but without requiring $\sum_{i=1}^{n}\alpha_{i}=1$. Checking these conditions can be formulated as linear programs with $n$ variables and $n+d+1$ or $n+d$ constraints respectively. These programs can be solved in a time polynomial in the dimension and the number of queries by interior-point methods, see e.g. Khachiyan (1979); Karmarkar (1984).
@@
To state the regret guarantees for CHC, we need to define the number $F(P)$ of flags of a polytope $P$. A flag of $P$ is a sequence $(F_{j})_{j=0}^{d-1}$ of faces [^1] Faces of a polytope are specific planar surfaces on its boundary, e.g., for a cube, the 0-dim faces are the vertices, the $1$-dim faces are the $12$ edges, and the $2$-dim faces are the $6$ squares forming the boundary. of $P$ such that $\dim(F_{j})=j$ and $F_{0}\subset\dots\subset F_{d-1}$. For example, for $d=2$ and $P=[0,1]^{2}$, a flag is a sequence formed by a vertex of the square and an edge incident to that vertex, and $F(P)=4\times 2=8$. Refer to Section 2.1 in Besau et al. (2018) for details. The following regret guarantees for CHC hold for any convex polytope partition $\{\mathcal{C}_{i}\}_{i\in[N]}$ of the embedding space by the expert.
@@
To upper bound $\mathbb{E}[\mu(\mathcal{C}_{i}\setminus\hat{\mathcal{C}}_{i,t})]$, we build upon a result from Bárány and Buchta (1993), which characterizes the expected volume of the convex hull of $n$ points sampled uniformly at random from a convex polytope in $\mathbb{R}^{d}$. We extend this result in two directions: (i) to distributions with bounded density via a rejection sampling argument; (ii) to the spherical setting where points lie on $\mathcal{S}^{d-1}$, by mapping spherical polytopes to Euclidean polytopes in $\mathbb{R}^{d-1}$ using the gnomonic projection. The dependence on $C/c$ (and $CK/c$ in the spherical case) arises from the rejection sampling step and is likely an artifact of the proof technique—we conjecture that this dependence could be removed with a tighter analysis.
@@
Voronoi Regret. We finally derive explicit regret upper bounds when the cells form a Voronoi tessellation, as described in Section 2.2. The following corollary is obtained from Theorem 3.1 by controlling the total number of flags, $\sum_{i=1}^{N}F(\mathcal{C}_{i})$. The proof, given in Appendix B.1.2, leverages an upper bound on the number of faces of a convex polytope from McMullen (1970), along with the observation that the number of facets of each cell $\mathcal{C}_{i}$ is at most N1+2dN-1+2d when $\mathcal{E}={\cal I}^{d}$ (the hypercube), and at most N1N-1 when $\mathcal{E}={\cal S}^{d-1}$ (the unit sphere).
@@
It may seem counterintuitive that the constant preceding the leading term (e.g., $\log^{d}T$ when ${\cal E}={\cal I}^{d}$) decreases with increasing $d$. This behavior suggests that the error term conceals a more intricate dependency on the dimension. Precisely characterizing this dependency is difficult and touches upon unresolved questions in convex geometry (see Appendix E.3 for further discussion). Nonetheless, our upper bounds on the regret of CHC are most informative in regimes where $T$ exceeds a threshold that grows exponentially with $d$, or equivalently, when the dimension is not excessively large. This is consistent with results from Chakraborti et al. (2021), which show that the expected volume of a convex hull formed by random points remains negligible until the number of samples is exponential in $d$. In our setting, this implies that a significant reduction in expert queries only begins once this sample threshold is crossed.